# Circle Squaring with Pieces of Small Boundary and Low Borel Complexity

### v3

@@ 5.3. Properties of the Constructed Sequences
Finally, in order to decide if some $\boldsymbol{v}\in{\mathbbm{T}}^{k}$ belongs to $I_{i}$, it is enough to compute the distance $d$ from $\boldsymbol{v}$ to $X_{i}$ in $G_{d}$ and then check if the $(d+5r^{\prime}_{i-1})$-neighbourhood of $\boldsymbol{v}$ contain a unique element of $X_{i}$. Since $d\leqslant r_{i}$, the set $I_{i}$ is a local function of $X_{i}$ of radius $r_{i}+5r^{\prime}_{i-1}\leqslant 2r_{i}$, finishing the proof of the lemma.
@@
Let us check the assumptions of Lemma 5.2. By induction, for every $1\leqslant j\leqslant i-1$, every component of $G_{d}\upharpoonright N_{q^{\prime}_{j-1}+2}[J_{j}]$ has diameter at most $(2r_{j}+2q^{\prime}_{j-1})+2(q^{\prime}_{j-1}+2)=t_{j}<t^{\prime}_{j}$ and the distance in $G_{d}$ between any two such components is at least $(5r^{\prime}_{j-1}-2q^{\prime}_{j-1})-2(q^{\prime}_{j-1}+2)$, which is strictly larger than $q^{\prime}_{j-1}+4$ by (5.8). Also, the distance between any two components of $G_{d}\upharpoonright I_{i}$ is at least $5r^{\prime}_{i-1}>q^{\prime}_{i-1}+4$ while the inequality in (5.3) holds (as equality). Thus Lemma 5.2 applies to the sequences in (5.14) and gives that $J_{i}\subseteq N_{q^{\prime}_{i-1}}[I_{i}]$, proving the first stated property.
@@
    $K_{i}\subseteq N_{q^{\prime}_{i-1}+2}[J_{i}]$, the distance in $G_{d}$ between any two distinct components of $G_{d}\upharpoonright K_{i}$ is least $5r_{i-1}^{\prime}-4q_{i-1}^{\prime}-4$, and every component $S\in\operatorname{comp}(K_{i})$ has diameter at most $t_{i}$ and contains exactly one vertex of $X_{i}$.
@@
    $L_{i}\subseteq N_{2r^{\prime}_{i}/5+q_{i}}[Y_{i}]$, the distance in $G_{d}$ between any two distinct components is least $r_{i}^{\prime}/5-2q_{i}$, and every component of $G_{d}\upharpoonright L_{i}$ has diameter at most $t^{\prime}_{i}$ and contains exactly one vertex of $Y_{i}$.
@@
Let us check the assumptions of Lemma 5.2 for the first triple of sequences. By induction, for every $1\leqslant j\leqslant i-1$, every component of $G_{d}\upharpoonright K_{j}$ has diameter at most $t_{j}$ and the distance in $G_{d}$ between any two such components is at least $5r^{\prime}_{j-1}-4q^{\prime}_{j-1}-4$, which is strictly more than $q^{\prime}_{j-1}+4$ by (5.8). Similarly, for each $1\leqslant j\leqslant i-2$, every component of $G_{d}\upharpoonright L_{j}$ has diameter at most $t^{\prime}_{j}$ and the distance in $G_{d}$ between any two such components is at least $r^{\prime}_{j}/5-2q_{j}$, which is strictly larger than $q_{j}+4$ by (5.9). Also, the inequality in (5.3) holds (as equality).
So, Lemma 5.2 applies and gives that $L_{i-1}\subseteq N_{q_{i-1}}[N_{2r^{\prime}_{i-1}/5}[Y_{i-1}]]$. This implies all other stated properties of $L_{i-1}$. Indeed, by Lemma 5.5, the distance between any two components of $G_{d}\upharpoonright L_{i-1}$, is at least $r^{\prime}_{i-1}/5-2q_{i-1}$, and each component has diameter at most $4r^{\prime}_{i-1}/5+2q_{i-1}\leqslant t^{\prime}_{i-1}$ and contains exactly one vertex from $Y_{i}$, proving all claims about $L_{i-1}$.
@@
Observe that, by Lemma 5.7, we have $S^{\prime}\subseteq K_{j}\subseteq N_{q^{\prime}_{j-1}+2}[J_{j}]$. Take the component $S$ of $G_{d}\upharpoonright J_{j}$ such that $S^{\prime\prime}:=N_{q^{\prime}_{j-1}+2}[S]$ intersects $S^{\prime}$. Let us show that, in fact, $S^{\prime}\subseteq S^{\prime\prime}$. Suppose on the contrary that we have some $\boldsymbol{u}\in S^{\prime}\setminus S^{\prime\prime}$. Again by Lemma 5.7, $\boldsymbol{u}$ is within distance $q^{\prime}_{j-1}+2$ to some component $X\in\operatorname{comp}(J_{j})$ that has to be different from $S$. But then the distance between $X$ and $S$ is, by the triangle inequality when we go via $S^{\prime}$, at most $2(q^{\prime}_{j-1}+2)+t_{i}<5r^{\prime}_{j-1}-2q^{\prime}_{j-1}$, contradicting Lemma 5.6.
@@
Let us check Property (2.17) for the sequence $(J_{i},K_{i})_{i=1}^{\infty}$. Note some asymmetry in (2.17), which states that if two components from two different times are at distance at most 2 then the 2-neighbourhood of the earlier one is contained in the later one. There are some cases to consider. First, take any $S\in\operatorname{comp}(J_{i})$. Suppose on the contrary to (2.17) that $S$ is at distance at most 2 from $X=J_{j}$ for $j>i$ or $X=K_{j}$ for $j\geqslant i$ but $N_{2}[S]$ is a not a subset of $X$. Let $T\in\operatorname{comp}(X)$ satisfy $\operatorname{dist}_{G_{d}}(S,T)\leqslant 2$. By construction, $T$ does not cut $N_{2}[S]$ as otherwise this set would be added into $T$ (since $T$ was build after $S$). Hence we have that $T\subseteq N_{2}[S]$. This is impossible for $j>i$ because the diameter of $S$ is at most $2r_{i}+2q_{i-1}^{\prime}$ by Lemma 5.6 while $T$ contains a ball of radius $\lfloor(r_{j}-5r_{i-1}^{\prime})/2\rfloor$ centred at the unique point of $X_{j}\cap T$ by the definition of $I_{j}\subseteq J_{j}\subseteq K_{j}$. Also, the case $X=K_{i}$ leads to a contradiction. Indeed, when constructing $T\in\operatorname{comp}(K_{i})$ we started with $N_{2}[T^{\prime}]$ for some $T^{\prime}\in\operatorname{comp}(J_{i})$, By Lemmas 5.4, 5.6 and 5.7, we must have $T^{\prime}=S$. Thus $K_{i}\supseteq T\supseteq N_{2}[T^{\prime}]$ contains $N_{2}[S]$, a contradiction. Second, take any $S\in\operatorname{comp}(K_{i})$. No set $J_{j}$ with $j>i$ can violate (2.17) by Lemma 5.8. Also, any $T\in\operatorname{comp}(K_{j})$ with $j>i$ and $\operatorname{dist}_{G_{d}}(S,T)\leqslant 2$ contains $N_{2}[S]$ (again since $T\subseteq N_{2}[S]$ is impossible by the diameter argument). Thus (2.17) holds for $(J_{i},K_{i})_{i=1}^{\infty}$.
@@
that is, we could have equivalently defined $\widetilde{K}_{i}$ by removing all individual vertices that belong to some $J_{j}$ with $j>i$. Likewise (with the claim following trivially from our definitions), we have
@@
For (5.20), take any $\boldsymbol{u}\in L_{i}\setminus\widetilde{L}_{i}$. It is enough to show that $\boldsymbol{u}\in\bigcup_{j=1}^{\infty}(J_{j}\cup\widetilde{K}_{j})$. By (5.18), there is $j>i$ with $\boldsymbol{u}\in K_{j}$. If $\boldsymbol{u}\in\widetilde{K}_{j}$ then we are done; otherwise the vertex $\boldsymbol{u}$ lies inside some $J_{h}$ with $h>j$ by (5.17), proving (5.20) in either case.
@@ 5.4. Some Structure of the Constructed Sets
For each $i\in{\mathbbm{N}}$, the set $J_{i}$ is a $4r_{i}$-local function of $X_{1},\dots,X_{i}$.
@@
We use induction on $i$ with the base case $i=1$ following from $J_{1}=I_{1}$ being in fact an $r_{1}$-local function of $X_{1}$. Let $i\geqslant 2$. In order to decide if some $\boldsymbol{v}\in{\mathbbm{T}}^{k}$ is included into $J_{i}$, it is enough to know all elements $\boldsymbol{u}\in X_{i}$ at distance at most $r_{i}$ from $\boldsymbol{v}$ and the component of $G_{d}\upharpoonright J_{i}$ containing each such $\boldsymbol{u}$. By Lemma 5.6, the component $S\in\operatorname{comp}(J_{i})$ containing $\boldsymbol{u}\in X_{i}$ has diameter at most $2r_{i}+2q^{\prime}_{i-1}$ and contains no other elements of $X_{i}$. So to build $S$ for given $\boldsymbol{u}\in X_{i}$, we need to know only the sets $N_{q^{\prime}_{0}+2}[J_{1}],\dots,N_{q^{\prime}_{i-2}+2}[J_{i-1}]$ inside $N_{2r_{i}+2q^{\prime}_{i-1}}[\boldsymbol{u}]$. By induction, each of these sets is a $4r_{i-1}$-local function of $X_{1},\dots,X_{i-1}$. The lemma follows since $r_{i}+(2r_{i}+2q^{\prime}_{i-1})+4r_{i-1}\leqslant 4r_{i}$.
@@
  • $\|\vec{n}-\vec{m}\|_{\infty}\leqslant r_{i}+5r^{\prime}_{i-1}$ and
@@
Suppose now that we split $\mathbbm{T}^{k}$ into Voronoi cells with the points in $X_{i}$ as the centres, similarly to the proof of Lemma 2.16. Each Voronoi cell contains $\Omega(r_{i}^{d})$ elements and, by the arguments above, $O(r_{i}^{d-1}r_{i-1}^{\prime})$ of these points are in $W_{i}$. Thus, the measure of $W_{i}$ is $O(r_{i-1}^{\prime}/r_{i})$.
@@ 5.5. Covering the Whole Torus
Now, for any $\boldsymbol{v}\in T_{\vec{p}}$, we have by definition of $R^{*}(\boldsymbol{v})$ that there are infinitely many indices $i$ for which there is a vertex $\boldsymbol{u}\in Y_{i}$ of the form $\boldsymbol{u}=\vec{n}\cdot_{a}\boldsymbol{v}$ for some $\vec{n}$ such that
@@ 6.1. Dealing with Finite Connected Sets
More generally, given a flow $\phi$ in $G_{d}$ and a set $D\subseteq\mathbbm{T}^{k}$ such that every set in $\operatorname{comp}(D)$ is finite, the operation of rounding $\phi$ along the boundary of $D$ is defined as follows. First, for every $S\in\operatorname{comp}(D)$ and every hole $S^{\prime}$ of $S$, we round $\phi$ along the boundary of $S^{\prime}$. Then, for each $S\in\operatorname{comp}(D)$, we round $\phi$ along the boundary of the union of $S$ and all of its holes. Each of these operations are well-defined and the order in which we perform them does not affect the result. Note that, for every finite connected $S$, the edge sets $\partial_{E}S^{\prime}$, for $S^{\prime}$ being a hole in $S$ or being $S$ with all its holes filled, partition the edge boundary of $S$. Indeed, every pair in $\partial_{E}S$ has exactly one point outside of $S$, lying in a (unique) component $S^{\prime}$ of $G_{d}\upharpoonright({\mathbbm{T}}^{k}\setminus S)$. If $S^{\prime}$ is not a hole of $S$, then $S^{\prime}$ is infinite while $S^{\prime\prime}$, the complement of $S^{\prime}$ in the component of $G_{d}$ containing $S$, is finite. Trivially, $S^{\prime\prime}$ cannot have any holes, so $S^{\prime\prime}$ is exactly $S$ with all its holes filled, as desired.
@@
Refer to captionRefer to caption

### v4

@@ 5.3. Properties of the Constructed Sequences
Finally, in order to decide if some $\boldsymbol{v}\in{\mathbbm{T}}^{k}$ belongs to $I_{i}$, it is enough to compute the distance $d$ from $\boldsymbol{v}$ to $X_{i}$ in $G_{d}$ and then check if the $(d+5r^{\prime}_{i-1}-1)$-neighbourhood of $\boldsymbol{v}$ contain a unique element of $X_{i}$. Since $d\leqslant r_{i}$, the set $I_{i}$ is a local function of $X_{i}$ of radius $r_{i}+5r^{\prime}_{i-1}\leqslant 2r_{i}$, finishing the proof of the lemma.
@@
Let us check the assumptions of Lemma 5.2. By induction, for every $1\leqslant j\leqslant i-1$, every component of $G_{d}\upharpoonright N_{q^{\prime}_{j-1}+2}[J_{j}]$ has diameter at most $(2r_{j}+2q^{\prime}_{j-1})+2(q^{\prime}_{j-1}+2)=t_{j}<t^{\prime}_{j}$ and the distance in $G_{d}$ between any two such components is at least $(5r^{\prime}_{j-1}-2q^{\prime}_{j-1})-2(q^{\prime}_{j-1}+2)$, which is strictly larger than $q^{\prime}_{j-1}+4$ by (5.8). Also, the distance between any two components of $G_{d}\upharpoonright I_{i}$ is at least $5r^{\prime}_{i-1}>q^{\prime}_{i-1}+4$ while the inequality in (5.3) for $1\leqslant j\leqslant i-1$ (which states that $q^{\prime}_{j}\geqslant t^{\prime}_{j}+2q^{\prime}_{j-1}+4$) holds by $q_{j}>q^{\prime}_{j-1}$. Thus Lemma 5.2 applies to the sequences in (5.14) and gives that $J_{i}\subseteq N_{q^{\prime}_{i-1}}[I_{i}]$, proving the first stated property.
@@
    $K_{i}\subseteq N_{q^{\prime}_{i-1}+2}[J_{i}]$, the distance in $G_{d}$ between any two distinct components of $G_{d}\upharpoonright K_{i}$ is least $5r_{i-1}^{\prime}-4q_{i-1}^{\prime}-4$, and every component of $G_{d}\upharpoonright K_{i}$ has diameter at most $t_{i}$ and contains exactly one vertex of $X_{i}$.
@@
    $L_{i}\subseteq N_{2r^{\prime}_{i}/5+q_{i}}[Y_{i}]$, the distance in $G_{d}$ between any two distinct components of $G_{d}\upharpoonright L_{i}$ is least $r_{i}^{\prime}/5-2q_{i}$, and every component of $G_{d}\upharpoonright L_{i}$ has diameter at most $t^{\prime}_{i}$ and contains exactly one vertex of $Y_{i}$.
@@
Let us check the assumptions of Lemma 5.2 for the first triple of sequences. By induction, for every $1\leqslant j\leqslant i-1$, every component of $G_{d}\upharpoonright K_{j}$ has diameter at most $t_{j}$ and the distance in $G_{d}$ between any two such components is at least $5r^{\prime}_{j-1}-4q^{\prime}_{j-1}-4$, which is strictly more than $q^{\prime}_{j-1}+4$ by (5.8). Similarly, for each $1\leqslant j\leqslant i-2$, every component of $G_{d}\upharpoonright L_{j}$ has diameter at most $t^{\prime}_{j}$ and the distance in $G_{d}$ between any two such components is at least $r^{\prime}_{j}/5-2q_{j}$, which is strictly larger than $q_{j}+4$ by (5.9). Also, the inequalities in (5.3) are routine to verify.
So, Lemma 5.2 applies and gives that $L_{i-1}\subseteq N_{q_{i-1}}[N_{2r^{\prime}_{i-1}/5}[Y_{i-1}]]$. This implies all other stated properties of $L_{i-1}$. Indeed, by Lemma 5.5, the distance between any two components of $G_{d}\upharpoonright L_{i-1}$, is at least $r^{\prime}_{i-1}/5-2q_{i-1}$, and each component has diameter at most $4r^{\prime}_{i-1}/5+2q_{i-1}=t^{\prime}_{i-1}$ and contains exactly one vertex from $Y_{i}$, proving all claims about $L_{i-1}$.
@@
Observe that, by Lemma 5.7, we have $S^{\prime}\subseteq K_{j}\subseteq N_{q^{\prime}_{j-1}+2}[J_{j}]$. Take a component $S$ of $G_{d}\upharpoonright J_{j}$ such that $S^{\prime\prime}:=N_{q^{\prime}_{j-1}+2}[S]$ intersects $S^{\prime}$. Let us show that, in fact, $S^{\prime}\subseteq S^{\prime\prime}$. Suppose on the contrary that we have some $\boldsymbol{u}\in S^{\prime}\setminus S^{\prime\prime}$. Again by Lemma 5.7, $\boldsymbol{u}$ is within distance $q^{\prime}_{j-1}+2$ to some component $X\in\operatorname{comp}(J_{j})$ that has to be different from $S$. But then the distance between $X$ and $S$ is, by the triangle inequality when we go via $S^{\prime}$, at most $2(q^{\prime}_{j-1}+2)+t_{i}<5r^{\prime}_{j-1}-2q^{\prime}_{j-1}$, contradicting Lemma 5.6.
@@
Let us check Property (2.17) for the sequence $(J_{i},K_{i})_{i=1}^{\infty}$. Note some asymmetry in (2.17), which states that if two components from two different times are at distance at most 2 then the 2-neighbourhood of the earlier one is contained in the later one. There are some cases to consider. First, take any $S\in\operatorname{comp}(J_{i})$. Suppose on the contrary to (2.17) that $S$ is at distance at most 2 from $X=J_{j}$ for $j>i$ or $X=K_{j}$ for $j\geqslant i$ but $N_{2}[S]$ is a not a subset of $X$. Let $T\in\operatorname{comp}(X)$ satisfy $\operatorname{dist}_{G_{d}}(S,T)\leqslant 2$. By construction, $T$ does not cut $N_{2}[S]$ as otherwise this set would be added into $T$ (since $T$ was built after $S$). Hence we have that $T\subseteq N_{2}[S]$. This is impossible for $j>i$ because the diameter of $S$ is at most $2r_{i}+2q_{i-1}^{\prime}$ by Lemma 5.6 while $T$ contains a ball of radius $\lfloor(r_{j}-5r_{i-1}^{\prime})/2\rfloor$ centred at the unique point of $X_{j}\cap T$ by the definition of $I_{j}\subseteq J_{j}\subseteq K_{j}$. Also, the case $X=K_{i}$ leads to a contradiction. Indeed, when constructing $T\in\operatorname{comp}(K_{i})$ we started with $N_{2}[T^{\prime}]$ for some $T^{\prime}\in\operatorname{comp}(J_{i})$, By Lemmas 5.4, 5.6 and 5.7, we must have $T^{\prime}=S$. Thus $K_{i}\supseteq T\supseteq N_{2}[T^{\prime}]$ contains $N_{2}[S]$, a contradiction. Second, take any $S\in\operatorname{comp}(K_{i})$. No set $J_{j}$ with $j>i$ can violate (2.17) by Lemma 5.8. Also, any $T\in\operatorname{comp}(K_{j})$ with $j>i$ and $\operatorname{dist}_{G_{d}}(S,T)\leqslant 2$ contains $N_{2}[S]$ (again since $T\subseteq N_{2}[S]$ is impossible by the diameter argument). Thus (2.17) holds for $(J_{i},K_{i})_{i=1}^{\infty}$.
@@
that is, we could have equivalently defined $\widetilde{K}_{i}$ from $K_{i}$ by removing all individual vertices that belong to some $J_{j}$ with $j>i$. Likewise (with the claim following trivially from our definitions), we have
@@
For (5.20), take any $\boldsymbol{u}\in L_{i}\setminus\widetilde{L}_{i}$. We have to show that $\boldsymbol{u}\in\bigcup_{j=1}^{\infty}(J_{j}\cup\widetilde{K}_{j})$. By (5.18), there is $j>i$ with $\boldsymbol{u}\in K_{j}$. If $\boldsymbol{u}\in\widetilde{K}_{j}$ then we are done; otherwise the vertex $\boldsymbol{u}$ lies inside some $J_{h}$ with $h>j$ by (5.17), proving (5.20) in either case.
@@ 5.4. Some Structure of the Constructed Sets
For each integer $i\geqslant 1$, the set $J_{i}$ is a $4r_{i}$-local function of $X_{1},\dots,X_{i}$.
@@
We use induction on $i$ with the base case $i=1$ following from $J_{1}=I_{1}$ being in fact an $(r_{1}+5r_{0})$-local function of $X_{1}$. Let $i\geqslant 2$. In order to decide if some $\boldsymbol{v}\in{\mathbbm{T}}^{k}$ is included into $J_{i}$, it is enough to know all elements $\boldsymbol{u}\in X_{i}$ at distance at most $r_{i}+q^{\prime}_{i-1}$ from $\boldsymbol{v}$ and the component of $G_{d}\upharpoonright J_{i}$ containing each such $\boldsymbol{u}$. By Lemma 5.6, the component $S\in\operatorname{comp}(J_{i})$ containing $\boldsymbol{u}\in X_{i}$ has diameter at most $2r_{i}+2q^{\prime}_{i-1}$ and contains no other elements of $X_{i}$. So to build $S$ for given $\boldsymbol{u}\in X_{i}$, we need to know only the sets $N_{q^{\prime}_{0}+2}[J_{1}],\dots,N_{q^{\prime}_{i-2}+2}[J_{i-1}]$ inside $N_{2r_{i}+2q^{\prime}_{i-1}}[\boldsymbol{u}]$. By induction, each of these sets is a $4r_{i-1}$-local function of $X_{1},\dots,X_{i-1}$. The lemma follows since $(r_{i}+q^{\prime}_{i-1})+(2r_{i}+2q^{\prime}_{i-1})+4r_{i-1}\leqslant 4r_{i}$.
@@
  • $\|\vec{n}-\vec{m}\|_{\infty}\leqslant r_{i}+5r^{\prime}_{i-1}$, and
@@
Suppose now that we split $\mathbbm{T}^{k}$ into Voronoi cells with the points in $X_{i}$ as the centres, similarly to the proof of Lemma 2.16. Each Voronoi cell contains $\Omega(r_{i}^{d})$ elements and, by the arguments above, $O(r_{i}^{d-1}r_{i-1}^{\prime})$ of these points are in $W_{i}$. It follows that the measure of $W_{i}$ is $O(r_{i-1}^{\prime}/r_{i})$.
@@ 5.5. Covering the Whole Torus
Now, for any $\boldsymbol{v}\in T_{\vec{p}}$, we have by the definition of $R^{*}(\boldsymbol{v})$ that there are infinitely many indices $i$ for which there is a vertex $\boldsymbol{u}\in Y_{i}$ of the form $\boldsymbol{u}=\vec{n}\cdot_{a}\boldsymbol{v}$ for some $\vec{n}$ such that
@@ 6.1. Dealing with Finite Connected Sets
More generally, given a flow $\phi$ in $G_{d}$ and a set $D\subseteq\mathbbm{T}^{k}$ such that every set in $\operatorname{comp}(D)$ is finite, the operation of rounding $\phi$ along the boundary of $D$ is defined as follows. First, for every $S\in\operatorname{comp}(D)$ and every hole $S^{\prime}$ of $S$, we round $\phi$ along the boundary of $S^{\prime}$. Then, for each $S\in\operatorname{comp}(D)$, we round $\phi$ along the boundary of the union of $S$ and all of its holes. All of these operations are well-defined and the order in which we perform them does not affect the result. Note that, for every finite connected $S$, the edge sets $\partial_{E}S^{\prime}$, for $S^{\prime}$ being a hole in $S$ or being $S$ with all its holes filled, partition the edge boundary of $S$. Indeed, every pair in $\partial_{E}S$ has exactly one point outside of $S$, lying in a (unique) component $S^{\prime}$ of $G_{d}\upharpoonright({\mathbbm{T}}^{k}\setminus S)$. If $S^{\prime}$ is not a hole of $S$, then $S^{\prime}$ is infinite while $S^{\prime\prime}$, the complement of $S^{\prime}$ in the component of $G_{d}$ containing $S$, is finite. Trivially, $S^{\prime\prime}$ cannot have any holes, so $S^{\prime\prime}$ is exactly $S$ with all its holes filled, as desired.
@@
Refer to captionRefer to captionRefer to caption